Check whether H={x∈ R*|x =1or x is irrational}and K={x∈ R*|x ≥1}are subgroups of (R*,.).


1
Expert's answer
2021-06-09T15:20:01-0400

Let us check whether "H=\{x\in R^*\ |\ x =1\text{ or }x\text{ is irrational }\}" is a subgroup of "(R^*,\cdot)". Since "\sqrt{2}\in H" and "\sqrt{2}\cdot\sqrt{2}=2\notin H", we conclude that "H" is not a subgroup of "(R^*,\cdot)".


Let us check whether "K=\{x\in R^*\ |\ x ≥1\}" is a subgroups of "(R^*,\cdot)". Taking onto account that for "x=2\in K" the inverse "x^{-1}=\frac{1}{2}\notin K," we conclude that "K" is not a subgroup of "(R^*,\cdot)".